A 0.5 F capacitor is connected to a battery pack with variable voltage. How would the energy stored in the capacitor change if the voltage across it doubled?

A. It wouldn't have any effect on the energy stored

B. The energy stored would double

C. The energy stored would halve

C. The energy stored would quadruple
Answer #1

capacitance (C) = 0.5 F

let voltage = V

so energy stored in capacitance (E) = 0.5( C) (V)²

now

new voltage (V') = 2V

so new energy ( E') = 0.5 (C) (2V)²

= 0.5 (C) (V)² (4)

E' = 4 E

C). THE ENERGY STORED WOULD QUADRUPLE.